OpenStudy (anonymous):

Evaluate \[\lim_{\theta \rightarrow 0} \frac{tan \theta}{\theta ^2 cot3\theta}\]

OpenStudy (anonymous):

use lim theta--->0 tan(theta)/theta=1

OpenStudy (anonymous):

Oh.. wait.. I thought I made some mistakes in previous calculation!

OpenStudy (anonymous):

lim theta--->0 tan(thata) tan (3theta)/ theta^2

OpenStudy (anonymous):

\[=\lim_{\theta \rightarrow 0} \frac{tan \theta}{\theta ^2cot3\theta}\]\[=\lim_{\theta \rightarrow 0} \frac{sin \theta}{\theta cos \theta}\frac{sin3\theta}{\theta cos3\theta}\]\[=(1)\lim_{\theta \rightarrow 0} 3\frac{sin3\theta}{3\theta cos3\theta}\]\[=(1)(3)=3\]

OpenStudy (anonymous):

That is it!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

*Note: \[\lim_{x \rightarrow 0}\frac{sinx}{x} = 1\]\[\lim_{x \rightarrow 0}\frac{sinx}{xcosx} = \lim_{x \rightarrow 0}\frac{sinx}{x} \times \lim_{x \rightarrow 0}\frac{1}{cosx} = 1 \]So, \[\lim_{x \rightarrow 0}\frac{tanx}{x}=1\]